OpenStudy (anonymous):

integrate X^1/2 sin^2 (X^3/2-1) dx

OpenStudy (anonymous):

substitute \[\LARGE u = x^{\frac{3}{2}} - 1\]\[\LARGE du = \frac{3}{2}x^{\frac{1}{2}}dx\]

OpenStudy (anonymous):

thanks, :)